### v1

# Big Reasoning with Small Models: Instruction Retrieval at Inference Time

Abstract
Can we bring big-scale reasoning to local-scale compute? Small language models (SLMs) are increasingly attractive for deployment because they run efficiently on local hardware, offering strong privacy, reduced cost, and lower environmental impact. Yet these models consistently struggle on tasks that require multi-step reasoning or specialized domain knowledge. To address this gap, we introduce instruction intervention at inference time. Our method builds an Instruction Corpus by grouping similar training questions and creating instructions that pair background knowledge with step-by-step procedures. At inference, the model retrieves the most relevant instructions and follows the steps. Unlike standard retrieval-augmented generation, which supplies raw passages, our approach provides structured scaffolds that guide reasoning. We evaluate this framework on MedQA (medical board exams), MMLU Professional Law, and MathQA using models from 3B to 14B parameters without any additional fine-tuning. Instruction retrieval yields consistent gains: 9.4% on MedQA, 7.9% on MMLU Law, and 5.1% on MathQA. Concise instructions outperform longer ones, and the magnitude of improvement depends strongly on model family and intrinsic reasoning ability.
Big Reasoning with Small Models: Instruction Retrieval at Inference Time
Kenan Alkiek, David Jurgens & Vinod Vydiswaran School of Information University of Michigan Ann Arbor, MI 48109, USA {kalkiek,jurgens,vgvinodv}@umich.edu
1 Introduction
Large language models (LLMs) such as GPT-4 have demonstrated remarkable generalization and reasoning ability across a wide range of domains, from mathematical problem solving (Cobbe et al., 2021) and clinical diagnostics (Kwon et al., 2024) to legal and commonsense reasoning (Hendrycks et al., 2021a). Their success stems from massive parameter scale and diverse pretraining data, which allow them to internalize both factual knowledge and multi-step reasoning procedures. However, scaling to hundreds of billions of parameters introduces substantial costs. For example, serving a single 175B-parameter model requires over 300GB of GPU memory and specialized infrastructure (Frantar et al., 2023; Zheng et al., 2022), making real-time or on-prem deployment difficult. Moreover, production-grade LLMs typically operate as closed-source APIs, raising privacy and governance concerns in sensitive domains such as medicine and law. Small language models offer a practical alternative. SLMs can be deployed locally, audited directly, and fine-tuned for specific domains (Pham et al., 2024; Belcak et al., 2025; Fu et al., 2023). Their compact size lowers cost and latency while enabling operation on commodity hardware, including laptops and edge devices. Moreover, recent work (Srivastava et al., 2025) challenges the notion that reasoning is exclusive to large-scale models and that smaller models can achieve competitive reasoning performance when trained or compressed effectively. Yet, their reasoning behavior remains underexplored Zhu et al. (2024), and unlike LLMs, which encode broad world knowledge and reasoning patterns within their parameters, SLMs still have a limited capacity to internalize such information Fu et al. (2023). This gap motivates the central question of this work: how can small, efficient models achieve strong reasoning performance without scaling up their parameters or relying on additional training?
Figure 1: Comparison of reasoning steps across model configurations on a MedQA-style case. The small language model alone produces an incorrect causal chain, while a large model reaches the right diagnosis but at high computational and privacy cost. Adding retrieved instructions supplies the missing background and procedural steps, enabling the SLM to reproduce the correct reasoning efficiently and privately.
We address this question through instruction intervention at inference time, which augments small models with explicit, retrievable reasoning procedures. Instead of generating reasoning steps it cannot reliably produce on its own, the model retrieves structured instructions that pair domain background with step-by-step guidance. As illustrated in Figure 1, this approach externalizes reasoning as a retrieval process, supplying the scaffolds that SLMs lack while retaining their efficiency and privacy advantages. The method is entirely inference-based: a single, unmodified SLM retrieves and follows instructions without additional training, and the approach generalizes across domains.
We evaluate this framework across three reasoning benchmarks (MedQA, MMLU Law, and MathQA) to assess both its effectiveness and the conditions under which it provides the greatest benefit. Specifically, we investigate two questions. First, does instruction retrieval reliably improve reasoning for small models, and what level of model capacity is needed for those gains to appear. Second, which properties of the retrieved instructions matter most. We test structure, length, and level of detail. We show that instruction retrieval improves accuracy by 5–10 percentage points for models above 3B parameters without additional fine-tuning. Concise, well-structured instructions that integrate factual knowledge with step-by-step reasoning yield the largest gains, and performance remains stable across clustering granularities, showing that the approach is robust to corpus design. Mixed-effects analysis indicates that improvement depends more on a model’s ability to follow structured guidance than on size. Notably, on knowledge-intensive tasks such as MedQA and MMLU Law, a 14B-parameter model equipped with retrieved instructions surpasses GPT-4o in zero-shot accuracy. Our results suggest that externalizing reasoning as retrievable text offers a practical path to scaling reliable inference on resource-limited or privacy-sensitive hardware. All code and the Instruction Corpus will be made publicly available upon acceptance to support reproducibility.
2 Related Work
Improving the reasoning ability of small language models has been approached from multiple directions. Chain-of-Thought (CoT) prompting encourages models to decompose a reasoning task into a sequence of intermediate steps rather than attempting to produce an answer directly. This approach has been shown to substantially improve the performance of LLMs across commonsense, symbolic, and mathematical reasoning benchmarks (Wei et al., 2023; Kojima et al., 2023; Wang et al., 2023). These gains, however, rely heavily on scale: only models with tens or hundreds of billions of parameters such as PaLM 540B (Chowdhery et al., 2022) or GPT-3 175B (Brown et al., 2020) reliably benefit from CoT prompting. Smaller models frequently generate illogical reasoning traces, and their accuracy can even decline when forced to produce step-by-step rationales (Wei et al., 2023).
Variant Avg. Length Know. Comp. Know. Rel. Sample Snippet (MedQA)
High School Concise 453 4.31 4.91 …See the big clue (underweight athlete with missed periods), match it to the rule (FHA $\rightarrow$ low estrogen), then pick “low bone density”…
Graduate Concise 631 4.57 4.92 …Prioritize the diagnostic triad (amenorrhea + low BMI + training), apply HPO suppression logic, reject violated conditions (e.g., high TSH/estrogen), and select “decreased bone density”…
[…]
Graduate Verbose 1,765 4.97 4.69 …Synthesize the amenorrhea–energy-deficit pattern, map it via leptin/kisspeptin $\rightarrow$ GnRH suppression to hypoestrogenism, weigh alternative axes (thyroid, prolactin, PCOS) by contradiction, and justify “decreased bone density” as the compelled outcome…
Table 1: Comparison of four instruction variants differing in audience level and length. Columns report the average token length and GPT-based quality scores (five-point Likert) for knowledge comprehensiveness and relevance, with example excerpts from MedQA. Concise variants are 2–3× shorter while retaining high relevance, whereas verbose variants expand coverage at the cost of slightly lower relevance.
To overcome the scale dependence of prompting, a parallel line of work distills reasoning traces from LLMs into SLMs. These methods fine-tune smaller models on CoT-style rationales generated by larger models or derived from labeled data, with the goal of internalizing step-by-step reasoning skills (Ho et al., 2023; Li et al., 2022; Magister et al., 2023; Fu et al., 2023; Hsieh et al., 2023). Distillation has proven effective on arithmetic and symbolic reasoning benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b), where reasoning patterns can be learned from repeated structures. However, the outcomes are less promising for knowledge-intensive tasks where accurate rationales depend on factual knowledge that smaller models struggle to retain. Because SLMs have limited parameter capacity, attempts to train them with extensive domain knowledge often lead to overspecialization and reduced generality (Fu et al., 2023).
Rather than encoding all knowledge directly into model parameters, retrieval-based methods expand a model’s capacity by supplementing it with external information at inference time. Classical dense retrieval (Karpukhin et al., 2020) improves open-domain question answering by retrieving relevant passages, but the returned evidence is often noisy or unstructured, leaving the model to extract and organize reasoning steps on its own. More recent generate-then-retrieve frameworks address this limitation by synthesizing targeted context before retrieval, yielding cleaner and more relevant evidence for reasoning (Yu et al., 2023). In high-stakes domains such as medicine, curated knowledge bases can further improve reliability: retrieving from medical textbooks, for example, outperforms relying on generic sources like Wikipedia for clinical decision-making (Wang et al., 2025). While these approaches have been developed for LLMs, their lessons are even more pertinent for SLMs. Smaller models are more easily overloaded by long or noisy contexts and are less capable of abstracting structure from raw passages. As a result, the effectiveness of retrieval for SLMs depends on ensuring that the retrieved evidence is relevant and structured.
Recent work (Kang et al., 2023; Zhao et al., 2024) improves upon unstructured retrieval by explicitly coupling evidence with the reasoning process itself. Knowledge-Augmented Reasoning Distillation (Kang et al., 2023) grounds LLM-style rationales in retrieved passages and employs a reranker to prioritize rationale-supporting evidence. This enables smaller models to learn reasoning tied to external knowledge, but still requires task-specific fine-tuning and keeps the reasoning policy embedded in parameters. Probe then Retrieve and Reason (Zhao et al., 2024) takes this modularity further by separating the pipeline into two distilled SLMs: a probing model that identifies what knowledge is needed and formulates retrieval queries, and a reasoning model that constructs step-by-step rationales from the retrieved passages. Our approach builds on the same intuition that retrieval should be directly tied to reasoning, but differs in where the reasoning resides. Rather than embedding procedures into model parameters or asking models to transform raw evidence into structured steps, we externalize reasoning into a reusable corpus of instructions. Our method largely separates reasoning from the model altogether.
Figure 2: Accuracy of SLMs with and without instruction retrieval across three reasoning benchmarks. Each pair shows zero-shot accuracy (black) and performance with High-School Concise instructions (green = improvement, red = decline). The dashed blue line marks GPT-4o’s zero-shot accuracy. Instruction retrieval yields consistent gains once models exceed 3B parameters, especially on MedQA and MMLU (Law), where knowledge and procedural reasoning are most demanding. Appendix Table 5 reports the same results in tabular form for reference.
3 Instruction Corpus
Small language models often fail on complex inference because they lack both sufficient domain knowledge and the structured support needed for multi-step reasoning. To address this gap, we construct the Instruction Corpus, a collection of modular instructions. Each instruction has two parts: (i) background knowledge relevant to a problem type and (ii) step-by-step reasoning procedures for solving it. At inference, instructions are retrieved and included in the prompt, guiding reasoning without task-specific fine-tuning. For example, in MedQA, a group of training questions may ask about contraindications for anticoagulants. From this group, we derive an instruction that outlines how to check mechanisms of action, review contraindications, and compare relative risks.
3.1 Corpus Construction
We construct the Instruction Corpus in three stages: clustering, instruction generation, and retrieval. First, training examples from each benchmark are embedded with OpenAI’s text-embedding-3-large and grouped using agglomerative clustering with average linkage and cosine distance. The resulting dendrogram supports different levels of granularity, where lower thresholds yield broad categories and higher thresholds produce highly specific clusters; by default, we adopt the most fine-grained partition so that each cluster corresponds to a distinct reasoning skill. Next, for each cluster we generate a reusable instruction by prompting GPT-5 with standardized templates (D) and up to 5 question examples from the cluster. Each instruction contains two sections: background knowledge summarizing the relevant domain concepts and a step-by-step procedure for solving new problems. The input examples serve only to guide generation and do not appear verbatim in the final instructions. Finally, at inference time, test queries are embedded with the same encoder and matched by cosine similarity to the closest clusters; the top-$k$ instructions (default $k=5$) are retrieved and included in prompt to provide both factual grounding and procedural scaffolding. See Appendix 2 for threshold values and statistics.
Figure 3: Accuracy differences when enforcing specific instruction styles compared to baseline instructions with no style or length constraints. Concise instructions generally outperform the baseline, while verbose instructions tend to reduce accuracy. Audience level (High School vs. Graduate) has smaller and less consistent effects.
3.2 Instruction Variants
While instructions provide missing knowledge and reasoning scaffolds, their effectiveness depends on how they are written. We therefore study how variation in audience level and length shapes small-model performance. For the audience level, we create two variants. A graduate-level version preserves domain terminology and assumes advanced conceptual familiarity (e.g., clinical reasoning in MedQA). The high-school-level version simplifies vocabulary and logical complexity (e.g. explain this to me like a high schooler). Aligning readability with prior knowledge reduces intrinsic cognitive load and improves comprehension, a core prediction of Cognitive Load Theory (Renkl & Atkinson, 2003; Sweller, 2024). Along the depth axis, we contrast concise instructions that contain only the essential reasoning steps with longer instructions that include additional commentary to stimulate germane load and schema formation. Moreover, recent prompt-engineering studies show that prompt compression often preserves, and sometimes even enhances, downstream accuracy despite large token reductions, which underscores the practical value of concise variants (Renze & Guven, 2024; Li et al., 2024). Crossing the two factors yields four prompt variants per task. Descriptive quality summaries by style are reported in Appendix Table 4. We also include a baseline prompt that does not enforce any constraints on style or length. This version reflects the instructions produced directly from our generation pipeline and serves as a neutral reference point. Comparing the four controlled variants against this baseline help us to isolate the effects of readability and verbosity from artifacts of the generation process. Full prompt templates for all conditions, including the baseline, are provided in Appendix D.
3.3 Quality Validation
Not all instructions are equally useful. Generic prompts such as “think carefully” may occasionally nudge performance but lack the structured knowledge and procedures that small models require. By contrast, the documents in our corpus are designed to provide both domain background and explicit reasoning steps. We verify that instructions are high quality and that the four variants differ only in style and length. To do so, we evaluate each instruction on three axes: knowledge (coverage and relevance of background facts), reasoning (soundness and task-specificity of steps), and clarity (structured, unambiguous presentation). A detailed rubric is given in Appendix A.1.3. Each document is scored automatically by GPT-5 on a five-point Likert scale. We repeat evaluations three times and report averaged scores; full results are shown in Table 4. Across tasks, mean scores exceed 4.7 on all three axes, confirming that the corpus provides accurate and reliable scaffolds rather than noisy or generic text. Table  1 summarizes quality differences across variants: verbosity increases comprehensiveness but slightly reduces relevance, while audience level has only minor effects. Clarity remains uniformly high across all conditions. Thus, the variants are equivalent in structure and quality, differing only in the controlled stylistic factors of length and audience.
3.4 Corpus Profile
The Instruction Corpus varies systematically in both length and size across tasks. Concise instructions average 500–750 tokens, while verbose instructions expand to 1.3k–2.1k tokens, a two- to threefold increase in prompt length. Length distributions are shown in Appendix Figure 7, with per-task breakdowns in Figure 8. Graduate-level instructions are slightly longer and more detailed than high-school versions, but the influence of audience framing is minor compared to verbosity. At the default, fine-grained clustering threshold, the corpus contains 29k groups for MathQA, 8.4k for MedQA, and 1.0k for MMLU Law. Group sizes are highly skewed: often reducing to single examples, while the largest contain up to 12. Each produces a narrowly scoped instruction often grounded in only one or a few examples, highlighting the specific cues and reasoning steps needed for that problem type. Broader thresholds merge these singletons into larger groups, yielding more general instructions. We return to this question in Section 6, where we show that instruction retrieval remains robust across a wide range of cluster sizes. Full statistics and length distributions are reported in Appendix 2 and Appendix A.1.1.
4 Experimental Setup
We evaluate the effect of the Instruction corpus on three benchmarks and publicly available models spanning five major open families, and compare performance against GPT-4o as a reference frontier model.
4.1 Tasks
We use three benchmarks that collectively stress different dimensions of reasoning. MedQA (Jin et al., 2020) contains multiple-choice questions from professional medical board exams; success requires both factual recall of biomedical knowledge and diagnostic reasoning. MMLU Professional Law (Hendrycks et al., 2021a) focuses on legal exam questions, testing recall of statutes and precedents alongside structured case-based reasoning. MathQA (Amini et al., 2019) evaluates symbolic reasoning through math word problems, requiring models to translate natural language into computational procedures.
4.2 Models
We evaluate across thirteen models from five major open families: Llama 3 (AI@Meta, 2024), Gemma 2 (Team, 2024), Qwen 3 (Team, 2025), Mistral (Jiang et al., 2023), and DeepSeek R1 Distilled (DeepSeek-AI, 2025). Model sizes span from 1B to 14B parameters, providing coverage across the typical range of deployable SLMs. For Llama 3, Gemma 2, and Mistral we use publicly released instruction-tuned variants. For DeepSeek R1 we evaluate distilled models released in sizes from 1.5B to 14B. For Qwen 3, we include both base models (1.7B, 4B, 14B) and an instruction-tuned variant (4B-Instruct). This mixture allows us to compare instruction-following capabilities directly against base model behavior. No additional fine-tuning is applied. At inference, each test question is embedded using the same model employed during clustering (text-embedding-3-large) and matched to training clusters by cosine similarity. The top five cluster-level instructions are retrieved and prepended to the query before decoding. To ensure comparability, all prompts and inference settings were standardized across tasks, with generation performed at a temperature of 0.7 and a top-p value of 0.95. GPT-4o, accessed via API, is included as a reference point to situate results against current frontier performance.
5 Results
Instruction retrieval reliably improves the reasoning ability of small language models across domains and architectures. Gains emerge once models reach a minimum capacity of around 3B parameters and grow with scale, ranging from 5 to 18 percentage points over zero-shot prompting. Improvements are largest on knowledge-intensive tasks such as MedQA and MMLU Law. To understand what drives these gains, we next examine how instruction design, model family, and prompt length shape performance, beginning with aggregate accuracy across tasks.
Figure 4: Marginal effects on accuracy relative to zero-shot prompting. All instruction variants yield large gains (+28–29pp). Family effects dominate: Qwen3 (+56pp) and DeepSeek R1 (+38–43pp) show the strongest ability to follow instructions compared to the LLaMA-3 reference, while Gemma-2 (+16pp) and Mistral (+2pp) are smaller. Model scale adds a consistent +9pp per log-unit increase in parameters, while longer prompts reduce accuracy by –8pp per log-token.
5.1 Aggregate Performance
Figure 2 shows the accuracy of the High School Concise instructions compared to zero-shot prompting across MathQA, MedQA, and MMLU Law. For the smallest models under 3B parameters, improvements are mixed and often negative. A few exceptions exist, such as Qwen-3 1.7B base, which gains +15pp on MedQA, but the same model shows negligible changes on MMLU and MathQA. This instability suggests that tiny models lack the intrinsic reasoning ability to reliably leverage retrieved scaffolds, which is in line with prior research Li et al. (2025). At around 3B parameters, gains become consistently positive, even if modest. LLaMA-3 3B, for instance, improves by +1–5pp across benchmarks. This marks a capacity threshold: once a model is large enough to follow multi-step guidance, retrieved instructions shift from being noise to providing usable structure. Above this threshold, performance consistently improves. Models from 4B to 14B shows gains across all three tasks, with deltas typically in the +5–18pp range. The largest improvements appear in MedQA and MMLU Law, where domain-specific knowledge and reasoning may lack representation in the limited parameter space. For example, Gemma-2 9B improves by +14pp on MMLU Law and Mistral 7B by +16pp, while MedQA gains reach +18pp at 9B. In contrast, MathQA gains are more muted, though still positive, likely because mathematical reasoning is already well represented in pretraining data, leaving less headroom for improvement. Overall, retrieved instructions consistently boost accuracy once models exceed 3B parameters, with robust gains across both symbolic and domain-expert reasoning tasks. Notably, on MedQA and MMLU Law, the 14B parameter SLMs with retrieved instructions surpass the zero-shot accuracy of GPT-4o.
5.2 Effects of Instruction Variation
Here, we compare the four instruction variations against a baseline condition in which models receive retrieved instructions without any explicit style or length constraints to learn more about the effects of how instructions are written, rather than the content of the instructions themselves. Figure 3 shows accuracy deltas across all variants. A consistent pattern is that concise instructions outperform the baseline, while verbose instructions often reduce accuracy, particularly for larger models. This finding mirrors our instruction-quality analysis: longer prompts increase comprehensiveness but dilute relevance, creating unnecessary cognitive load for models that are already capable of multi-step reasoning.
Audience effects are smaller and less consistent. In some cases, simplified high school variants provide a slight benefit, especially on MedQA, where complex biomedical terminology may overwhelm smaller models. In other cases, graduate-level detail performs equally well or marginally better. Overall, these results suggest that the primary determinant of instruction effectiveness is length rather than audience level. Concise scaffolds reduce context overhead while still supplying the essential reasoning steps, aligning with our quality evaluation that verbosity may introduce tangential material. The negligible effect of audience framing implies that once instructions are well structured, SLMs can adapt to register differences without measurable loss. Indicating that instruction retrieval benefits from minimizing extraneous detail rather than tailoring explanations to a presumed audience.
5.3 Determinants of Instruction Effectiveness
To isolate which properties of instructions and models drive performance gain we fit a mixed-effects logistic regression at the question–model–variant level (Figure 4). The dependent variable is a binary indicator of whether the model answered a test question correctly and fixed effects include instruction variant (audience and length), prompt length, model size, and a zero-shot indicator. Interactions with model size test whether stylistic effects vary with capacity. Random intercepts for questions and datasets account for variation in item difficulty and domain. LLaMA-3 serves as the reference family, chosen because it is the oldest model in our pool. Coefficients are reported as average marginal effects, which represent the change in predicted accuracy when a predictor increases by one unit while other variables are held constant. Positive values indicate improvements relative to the reference categories.
Both High School and Graduate instruction variants yield large and significant gains of +28–29pp over zero-shot prompting, confirming that retrieved scaffolds substantially improve small-model reasoning. Yet among fixed effects, model family dominates: Qwen-3 shows the largest positive offset (+56pp), followed by DeepSeek R1–LLaMA (+43pp) and DeepSeek R1–Qwen (+38pp). These gains are substantially larger than the +9pp associated with a log-unit increase in size, highlighting that architecture and pretraining choices matter as much as, or more than, scale. Families like Qwen and DeepSeek R1 may already have stronger step-following and reasoning performance out of the box compared to the older Llama3 variants. Even with the same parameter count, families differ in attention implementations, training recipes, or RLHF alignment, all of which can influence how reliably they follow structured prompts. Prompt length has a negative effect. Each log-unit increase in tokens reduces accuracy by –8pp, highlighting that verbosity imposes a systematic penalty. While conciseness does not improve accuracy on average, its interaction with size is positive: larger models benefit disproportionately from brevity, while smaller models gain little. Verbose instructions add no measurable value, again suggesting that detail beyond essential steps creates redundancy rather than a usable signal.
6 Ablation and Analysis
While instruction retrieval improves small-model reasoning, the source of these gains remains unclear. We perform targeted ablations on MedQA using the High School Concise variant and models of at least 3B parameters to identify what drives improvement.
6.1 Component Ablations
Do instruction gains stem from retrieved background knowledge, reasoning steps, or their combination? To isolate these effects, we ablate the two components of each instruction on MedQA. In the knowledge-only condition, instructions include only the background section, providing factual context but no reasoning steps. In the reasoning-only condition, instructions include only the procedural steps, with no supporting knowledge. The full condition retains both.
Figure 5: Performance impact of ablating instruction components on MedQA (High School Concise). Both knowledge-only and reasoning-only variants outperform zero-shot prompting, though the relative advantage differs by model family.
Both components independently improve performance over the zero-shot baseline (Figure 5). On average, removing either section reduces accuracy relative to the full instruction, confirming that background knowledge and procedural steps contribute complementary information. Across models, the knowledge-only variant outperforms the reasoning-only variant by roughly two percentage points on average, suggesting that factual grounding plays a slightly stronger role in MedQA. The pattern, however, varies by model family: Qwen and DeepSeek models benefit most from factual context, whereas Gemma gains more from reasoning steps. In short, impact varies with each model’s strengths: models lacking domain knowledge rely more on factual context, whereas those with limited reasoning ability benefit more from explicit steps.
6.2 Cluster Size
The instruction corpus groups training questions into clusters using agglomerative clustering with cosine distance thresholds (Appendix Table 2). Each clustering threshold determines how close examples must be to form a cluster, with lower values producing many small, fine-grained clusters and higher values merging examples into broader, more general groups. From a design perspective, smaller clusters may yield highly specific instructions that fit narrow problem types but limit reuse, while larger clusters may promote generalization at the risk of losing task-specific cues. We therefore evaluate how this granularity influences downstream accuracy. Figure 6 reports average MedQA accuracy across models as a function of mean cluster size. Accuracy remains stable as clusters become broader, indicating that instructions can represent larger groups of questions without loss in performance. This robustness suggests that clustering granularity can be treated as a tunable hyperparameter: by adjusting the threshold on a development set, we can identify the coarsest grouping that preserves accuracy. Optimizing this allows the use of fewer instructions that are more representative of topics rather than individual questions and may improve interpretability. Per-model accuracy across thresholds is reported in Appendix Table 6.
7 Discussion
Our findings show that small models above a minimal capacity threshold can reliably incorporate retrieved instructions to perform multi-step reasoning, narrowing the gap to large models without additional training. This pattern aligns with prior evidence (Li et al., 2025) that models under roughly 3B parameters struggle to internalize long reasoning traces but can effectively leverage shorter, externally provided guidance. By treating instructions as modular, retrievable assets, our approach shifts part of the reasoning process from model parameters to structured text, creating a practical foundation for more maintainable and auditable small models. Externalized instructions are non-parametric and can be updated or versioned without retraining, enabling faster and safer revisions in regulated domains. Because they are written in natural language, instructions are also interpretable and portable across model families, allowing a single high-quality corpus to provide consistent reasoning support for diverse architectures. In applied settings, this separation makes reasoning assistance feasible in privacy-constrained environments: for example, a 7B-parameter model running locally on a clinician’s laptop can retrieve targeted instructions for diagnostic reasoning or documentation without transmitting patient data. Ultimately, instruction effectiveness depends on a model’s existing competence and confidence. Domains where models already exhibit strong internal reasoning, such as mathematics, show smaller gains from retrieval because the external instructions provide limited additional value. The same principle likely applies at the instance level: even within a single model, some queries require external guidance while others can be solved unaided. Future systems should therefore adapt retrieval based on a model’s estimated capability and uncertainty, invoking instructions only when they are expected to improve reasoning. Developing such competence-aware or confidence-gated retrieval mechanisms is a key step toward making instruction intervention dependable.
Figure 6: Effect of clustering granularity on MedQA accuracy, averaged across all evaluated models. Performance remains consistent across thresholds, indicating that instruction retrieval is robust even when clusters merge into broader groups.
8 Conclusion
This work reframes reasoning as a retrieval problem: rather than generating reasoning chains from scratch, small language models retrieve structured instructions that integrate background knowledge with step-by-step reasoning to guide their generation. Across MedQA, MMLU Law, and MathQA, this approach consistently improves multi-step reasoning without additional training. Concise, well-structured instructions yield the largest gains, and improvement depends more on a model’s ability to follow structure than on size alone. These findings suggest that both factual knowledge and reasoning procedures can be externalized as reusable text modules, enabling small models to approximate large-model performance while preserving efficiency and privacy. Looking ahead, competence-aware retrieval mechanisms that adapt instruction use based on model confidence may further bridge the gap between compact and frontier-scale reasoning systems.
Limitations
Several limitations remain. The current corpus is derived from benchmark datasets rather than the open-ended problems encountered in real-world use. Extending this framework to domain-scale or continuously evolving environments raises a broader question: how should instructions and their associated knowledge be represented, retrieved, and maintained in the wild rather than within fixed benchmarks? Our retrieval quality also depends on a simple top-$k$ similarity search; future work could incorporate re-ranking or adaptive selection to improve relevance and coverage as the corpus grows.
Appendix A Appendix
A.1 Instruction Corpus
A.1.1 Length Distributions
Figure 7: Distribution of instruction lengths across all variants. Concise instructions average 500–750 tokens, while verbose versions expand to 1.3k–2.1k tokens, a consistent two- to threefold increase
Figure 8: Distribution of instruction lengths across all variants. and broken down by task.
A.1.2 Clustering Details
Task Threshold # Clusters Mean Size Std. Size Max Size Silhouette
MedQA 0.177 8414 1.21 0.63 11 0.079
0.209 7070 1.44 1.08 18 0.107
0.244 5390 1.89 1.83 20 0.127
0.272 4130 2.46 2.74 38 0.135
0.325 2198 4.63 6.34 62 0.128
MMLU 0.246 1048 1.30 0.79 12 0.083
0.275 872 1.57 1.29 21 0.101
0.322 586 2.33 2.67 36 0.101
0.352 410 3.33 4.45 49 0.099
0.635 3 455.7 639.5 1360 0.123
MathQA 0.008 29093 1.03 0.22 7 0.024
0.063 15949 1.87 1.88 30 0.387
0.093 12229 2.44 2.81 61 0.480
0.154 8509 3.51 4.97 140 0.539
0.255 4789 6.23 6.23 786 0.442
Table 2: Cluster statistics across thresholds for MedQA, MMLU (Law), and MathQA. We report the number of clusters, mean and standard deviation of cluster size, maximum cluster size, and silhouette score (higher = more cohesive clusters).
A.1.3 Evaluation Criteria
Criterion Score Descriptor
Knowledge Comprehensiveness 5 All background facts needed for typical instances are present, including key definitions, edge cases, and disambiguations.
4 Nearly all essentials covered. Minor omissions that rarely affect correctness.
3 Mixed coverage with several common cases or definitions missing. Sometimes blocks a correct solution.
2 Multiple essential facts missing. Frequent failure without extra knowledge.
1 Largely incomplete background.
Knowledge Relevance 5 Background contains only necessary or highly useful facts for the cluster. No tangents.
4 Small amount of extra detail that is not distracting.
3 Noticeable extraneous content. Can distract or slow reasoning.
2 Large amount of irrelevant or low-yield content. Likely to mislead.
1 Mostly off-topic or generic background.
Reasoning Accuracy 5 Reasoning steps are logically sound, factually correct, and properly sequenced. No contradictions.
4 Minor imprecision or wording issues that do not change the outcome.
3 At least one underspecified or brittle step that could lead to a wrong branch. Small gaps.
2 Clear logical or factual error that would often yield an incorrect answer.
1 Reasoning is largely incorrect or inconsistent.
Reasoning Relevance 5 Steps are tailored to the problem type, align with input and output, and include cluster-critical operations.
4 Mostly tailored with minimal generic filler.
3 Mix of tailored and generic steps. Some do not map to the task structure.
2 Largely generic advice. Missing one or more cluster-critical steps.
1 Steps unrelated to the problem type.
Clarity 5 Concise, well structured, unambiguous. Consistent terminology. Numbered steps or clear bullets. Explicit stop conditions or decision points.
4 Generally clear with minor verbosity or mild ambiguity.
3 Mixed clarity. Some steps vague or terminology inconsistent.
2 Hard to follow. Long sentences, unclear step boundaries, or undefined terms.
1 Confusing or unreadable.
Table 3: Five-point Likert rubric for instruction quality. Global decision rules: cap Reasoning Accuracy at 2 if any factual error is present in the steps. Cap Reasoning Relevance at 2 and Knowledge Comprehensiveness at 3 if a required step is missing. Cap Knowledge Relevance at 2 if much of the background is tangential. Cap Clarity at 3 if step boundaries are unclear or terminology is inconsistent.
A.1.4 Variant Evaluation Results
Audience Length Knowledge Comp. Knowledge Rel. Reasoning Acc. Reasoning Rel. Clarity
Graduate Concise 4.57 4.92 4.83 4.90 4.95
Graduate Long 4.97 4.69 4.93 4.98 4.94
High school Concise 4.31 4.91 4.73 4.81 4.91
High school Long 4.68 4.82 4.90 4.95 4.91
Table 4: Instruction quality scores (five-point Likert scale) across audience level and length variants.
Appendix B Results
B.1 Inference Settings
All models were evaluated using a consistent inference configuration. Generation employed a temperature of 0.7 and a top-p value of 0.95 across all tasks. For fairness, each model operated at its default context length (typically 8,192 tokens for smaller models). When the concatenated prompt and retrieved instructions exceeded the model’s context window, the number of retrieved instructions was reduced from the top-5 matches (e.g., using the top-4) until the full prompt fit within the allowable limit. This ensures uniform decoding behavior while maintaining maximal use of retrieved guidance within each model’s constraints.
B.2 Aggregate Performance Results
Task Model Instruction Acc. Zero-Shot Acc. $\Delta$
MathQA
LLaMA-3 1B Instruct 0.07 0.10 -0.03
DeepSeek-R1 Qwen 1.5B 0.78 0.68 +0.10
Qwen-3 1.7B Base 0.90 0.88 +0.01
Gemma-2 2B IT 0.36 0.39 -0.02
LLaMA-3 3B Instruct 0.15 0.11 +0.04
Qwen-3 4B Base 0.91 0.91 +0.01
Qwen-3 4B Instruct 0.91 0.90 +0.01
DeepSeek-R1 Qwen 7B 0.88 0.88 +0.00
Mistral 7B Instruct 0.35 0.22 +0.14
DeepSeek-R1 LLaMA 8B 0.87 0.84 +0.03
Gemma-2 9B IT 0.75 0.64 +0.10
DeepSeek-R1 Qwen 14B 0.90 0.89 +0.02
Qwen-3 14B Base 0.91 0.79 +0.12
MedQA
LLaMA-3 1B Instruct 0.19 0.30 -0.11
DeepSeek-R1 Qwen 1.5B 0.31 0.21 +0.10
Qwen-3 1.7B Base 0.60 0.45 +0.15
Gemma-2 2B IT 0.32 0.36 -0.04
LLaMA-3 3B Instruct 0.50 0.49 +0.01
Qwen-3 4B Base 0.73 0.64 +0.08
Qwen-3 4B Instruct 0.74 0.64 +0.10
DeepSeek-R1 Qwen 7B 0.47 0.34 +0.12
Mistral 7B Instruct 0.44 0.33 +0.11
DeepSeek-R1 LLaMA 8B 0.60 0.51 +0.09
Gemma-2 9B IT 0.54 0.37 +0.18
DeepSeek-R1 Qwen 14B 0.74 0.68 +0.06
Qwen-3 14B Base 0.83 0.73 +0.09
MMLU (Law)
LLaMA-3 1B Instruct 0.20 0.31 -0.11
DeepSeek-R1 Qwen 1.5B 0.29 0.31 -0.01
Qwen-3 1.7B Base 0.46 0.45 +0.01
Gemma-2 2B IT 0.28 0.30 -0.02
LLaMA-3 3B Instruct 0.38 0.34 +0.05
Qwen-3 4B Base 0.57 0.47 +0.11
Qwen-3 4B Instruct 0.54 0.49 +0.05
DeepSeek-R1 Qwen 7B 0.36 0.37 -0.01
Mistral 7B Instruct 0.38 0.22 +0.16
DeepSeek-R1 LLaMA 8B 0.53 0.49 +0.04
Gemma-2 9B IT 0.41 0.26 +0.14
DeepSeek-R1 Qwen 14B 0.61 0.55 +0.07
Qwen-3 14B Base 0.65 0.53 +0.12
Table 5: Aggregate performance for the High School Concise instruction variant. Instruction retrieval consistently improves performance across tasks and models once capacity exceeds 3B parameters, with the largest gains on MedQA and MMLU Law. $\Delta$ denotes the difference between instruction and zero-shot accuracy.
Appendix C Ablation
Model Threshold 1 Threshold 2 Threshold 3 Threshold 4 Threshold 5
Avg. Cluster Size
1.21 1.44 1.89 2.46 4.63
Avg. Accuracy (%)
deepseek_r1_llama_8b 62.0 61.1 59.1 59.6 57.6
deepseek_r1_qwen_14b 76.0 74.3 75.0 75.2 72.1
deepseek_r1_qwen_7b 45.3 45.0 40.8 38.7 35.8
gemma_2_9b_it 47.8 45.8 47.1 46.1 46.8
llama3_3b_instruct 46.1 47.1 49.3 48.4 46.6
mistral_7b_instruct 42.2 45.0 46.8 45.6 47.1
qwen3_14b_base 81.6 82.8 82.7 82.4 81.8
qwen3_4b_base 73.8 74.5 74.5 72.3 72.2
qwen3_4b_instruct 75.1 75.5 73.9 75.3 73.3
Table 6: Performance on MedQA across clustering thresholds by model. Each column corresponds to a clustering threshold, showing the average cluster size and model accuracy (%).
Appendix D Instruction Prompt Templates
You are tasked with creating an instruction guide based on a set of examples of similar problems. Carefully examine these examples to identify common patterns, concepts, and problem-solving approaches. Your analysis should focus on extracting knowledge and reasoning patterns.
Consider the examples as a whole to understand their complexity and domain. Let the nature of the questions guide how simple or sophisticated and detailed your instructions should be. Write in clear, instructional language as if teaching someone how to solve this type of problem, using terminology and explanations that match the level demonstrated in the examples.
Your response must contain exactly these two sections with these exact headers:
## Background Knowledge
Present the essential principles, definitions, or rules that are directly relevant across the examples. Include important patterns that help distinguish correct from incorrect answers. For each key point, indicate whether it strongly determines the answer when present, provides helpful support, or might mislead if given too much weight. Use the language and concepts appropriate to the field and complexity level shown in the examples.
## Reasoning Steps
Provide a clear approach that works across examples, connecting each step to the background knowledge. Start by identifying what the question is asking and what key information or clues to look for. Explain how to apply the most important knowledge first, and when certain clues should override other considerations. Address how to weigh different types of evidence and handle situations where answers might seem similar. When relevant, explain how to tell apart commonly confused options. End by stating what should determine the final choice.
Here are the examples to analyze:
<examples>
{EXAMPLES}
</examples>
The instructions must work for every example provided.
Figure 9: Baseline Prompt Template
You are tasked with creating a **quick, practical** instruction guide for **high school students** based on a set of examples of similar problems. Carefully examine these examples to identify common patterns, concepts, and problem-solving approaches. Your analysis should focus on extracting knowledge and reasoning patterns.
Write in **clear, simple language** that high school students can understand quickly. Keep your response **short and practical** - focus only on what students need to know to solve the problem. Your response must contain exactly these two sections with these exact headers:
## Background Knowledge
Present the essential concepts needed to solve these problems. Include important patterns that help distinguish correct from incorrect answers. For each key point, indicate whether it strongly determines the answer when present, provides helpful support, or might mislead if given too much weight. **Use simple language and keep explanations short and practical.**
## Reasoning Steps
Provide a clear approach that works across examples, connecting each step to the background knowledge. Start by identifying what the question is asking and what key information or clues to look for. Explain how to apply the most important knowledge first, and when certain clues should override other considerations. Address how to weigh different types of evidence and handle situations where answers might seem similar. When relevant, explain how to tell apart commonly confused options. End by stating what should determine the final choice. **Keep steps clear and practical.** Here are the examples to analyze:
<examples>
{EXAMPLES}
</examples>
The instruction guide must work for every example provided.
Figure 10: High School Concise Prompt Template
You are tasked with creating a **comprehensive, encouraging** instruction guide for **high school students** based on a set of examples of similar problems. Carefully examine these examples to identify common patterns, concepts, and problem-solving approaches. Your analysis should focus on extracting knowledge and reasoning patterns.
Write in **clear, encouraging language** that builds confidence for high school students. Provide **detailed explanations** to help students really understand. Your response must contain exactly these two sections with these exact headers:
## Background Knowledge
Present the essential concepts that directly help solve these problems. Include important patterns that help distinguish correct from incorrect answers. For each key point, indicate whether it strongly determines the answer when present, provides helpful support, or might mislead if given too much weight. **Use encouraging, detailed language with examples.** Connect new concepts to things students already know and explain why principles work the way they do.
## Reasoning Steps
Provide a detailed approach that works across examples, connecting each step to the background knowledge with encouraging explanations. Start by identifying what the question is asking and what key information or clues to look for. Explain how to apply the most important knowledge first, and when certain clues should override other considerations. Address how to weigh different types of evidence and handle situations where answers might seem similar. When relevant, explain how to tell apart commonly confused options by describing the key differences. End by stating what should determine the final choice and explain why. **Remember to explain the "why" behind each step to build understanding and confidence.** Here are the examples to analyze:
<examples>
{EXAMPLES}
</examples>
The instruction guide must work for every example provided.
Figure 11: High School Verbose Prompt Template
You are tasked with creating a **concise** instruction guide for **graduate-level students** based on a set of examples of similar problems. Carefully examine these examples to identify common patterns, concepts, and problem-solving approaches. Your analysis should focus on extracting knowledge and reasoning patterns.
Write in precise, **graduate-level language** using terminology appropriate for the domain. Keep your response **concise and focused on essential information only**. Your response must contain exactly these two sections with these exact headers:
## Background Knowledge
Present the essential principles, definitions, and domain knowledge directly relevant across most examples. Include important patterns that help distinguish correct from incorrect answers, even if they appear less frequently. For each key point, indicate whether it strongly determines the answer when present, provides helpful support, or might mislead if given too much weight. **Use precise graduate-level terminology.**
## Reasoning Steps
Provide a systematic approach that works across examples, connecting each step to the background knowledge. Start by identifying what the question is asking and what key information or clues to look for. Explain how to apply the most important knowledge first, and when certain clues should override other considerations. Address how to weigh different types of evidence and handle situations where answers might seem similar. When relevant, explain how to tell apart commonly confused options. End by stating what should determine the final choice. **Keep explanations concise but complete.** Here are the examples to analyze:
<examples>
{EXAMPLES}
</examples>
The instruction guide must work for every example provided.
Figure 12: Graduate Concise Prompt Template
You are tasked with creating a **comprehensive, thorough** instruction guide for **graduate students** based on a set of examples of similar problems. Carefully examine these examples to identify common patterns, concepts, and problem-solving approaches. Your analysis should focus on extracting knowledge and reasoning patterns.
Write in precise, **graduate-level language** using **advanced terminology** appropriate for the domain. Provide **comprehensive coverage and detailed explanations**. Your response must contain exactly these two sections with these exact headers:
## Background Knowledge
Present comprehensive coverage of the essential principles, definitions, and domain knowledge directly relevant across most examples. Include important patterns that help distinguish correct from incorrect answers, even if they appear less frequently. For each key point, indicate whether it strongly determines the answer when present, provides helpful support, or might mislead if given too much weight. **Use precise graduate-level terminology.** Include detailed definitions with theoretical context, fundamental principles, complex relationships between concepts, and connections to broader theory when directly applicable.
## Reasoning Steps
Provide a comprehensive approach that works across examples, connecting each step to the background knowledge with detailed explanations. Start by identifying what the question is asking and what key information or clues to look for. Explain how to apply the most important knowledge first, and when certain clues should override other considerations. Address how to weigh different types of evidence and handle situations where answers might seem similar, explaining the theoretical foundations behind this hierarchy. When relevant, explain how to tell apart commonly confused options by comprehensively comparing the alternatives. Include multiple solution approaches when applicable, discuss trade-offs between methods, and address edge cases. End by stating what should determine the final choice with theoretical justification. Here are the examples to analyze:
<examples>
{EXAMPLES}
</examples>
The instruction guide must work for every example provided.
Figure 13: Graduate Verbose Prompt Template